# Neil A. Chriss

### v1

@@ 3.6 Bayesian equilibrium
Therefore we can alternatively define $M:=-(A^{T}A)^{-1}A^{T}B$ in the case where $A$ may be singular. In the case $A$ is non-singular this is equivalent to 3.19 because as is well-known $(A^{T}A)^{-1}A^{T}=A^{-1}$ (when $A$ is non-singular ) [^10] This is because when $A$ is non-singular then $(A^{T}A)^{-1}=A^{-1}(A^{T})^{-1}$ then $(A^{T}A)^{-1}A^{T}=A^{-1}(A^{T})^{-1}A^{T}=A^{-1}$.. In the next section we use this to explicitly solve for the relevant strategies. In general we will proceed without reference to the normal-form equations but note that these should be employed when $A$ may not singular or nearly singular.
@@
We set our aim to find the most straightforward description of the matrix $C$. To begin we construct matrices $\Pi_{1}$, $\Pi_{2}$ from the sub-blocks of equations (3.12) and (3.13) consisting of conditional probabilities:
@@
$M=-A^{-1}B$ $=\begin{pmatrix}-(I-\Pi_{1}\Pi_{2})^{-1}&\Pi_{1}(I-\Pi_{2}\Pi_{1})^{-1}\\ \Pi_{2}(I-\Pi_{1}\Pi_{2})^{-1}&-(I-\Pi_{2}\Pi_{1})^{-1}\end{pmatrix}\begin{pmatrix}0&D_{1}\Pi_{1}\\ D_{2}\Pi_{2}&0\end{pmatrix}$
@@
Note that Eq. 3.33 provides a complete description of a two-firm Bayesian equilibrium of the two normalized strategies: $s_{1}^{*}=(s_{1}^{1},\dots,s_{1}^{K})$ and $s_{2}^{*}=(s_{2}^{1},\dots,s_{2}^{M})$ satisfying the boundary conditions in Eq. 3.31. We also note that from Eq. 4.2b we immediately have:
@@ 3.7 Two firms competing in the presence a non-strategic firm
When two-firms are in competition with one another and there is the possibility of a third firm trading non-strategically, that is, without taking into account how its competition is trading, we may analyze the situation using a modified version of the tools we have already developed. In this situation let us assume that neither firm is sure whether or not there is actually a third firm. Thus in terms of types (see Section 2.3) it may be that one firm’s particular type believes there is indeed a non-strategic firm trading a certain target quantity, while another type believes there is no such firm. This section presents Bayesian equilibria for this situation. But first we make a small historical comment.
@@
    We will write $K_{1}$ and $K_{2}$ for the number of types for firm 1 and 2 respectively.
@@
    Each possible type of firms 1 and 2 believes that the non-strategic firm will trade a risk-neutral strategy with a specific target size. We write $b_{i}^{k}$ for firm $i$, type $k$’s belief about the target size of the non-strategic firm. Note that it is possible that some of the $b_{i}^{k}$ are equal to zero meaning that firm $i$, type $k$ believes there not a non-strategic investor.
@@
    firms 1 and 2 have $K_{1}$ and $K_{2}$ types respectively and the type data for firm $i$, type $k$ is now a triple $(\kappa_{i}^{k},f_{i}^{k},b_{i}^{k})$, where $\kappa_{i}^{k}$ is the market impact parameter, $f_{i}^{k}$ is the target size and $b_{i}^{k}$ is the target size that firm $i$, type $k$ believes the non-strategic firm will trade.
@@
    There is a common prior probability matrix $P$, see Eq. 2.3, such that $p_{km}$ is the join probability that firm 1’s active type is $k$ and firm 2’s is $m$.
@@
We note that it is possible and straightforward to extend the analysis we do so that some firms types believe that the non-strategic firm trades some other strategy other than the risk-neutral strategy and our analysis directly extends to this case so long as the other other strategy’s functional form is fixed and known ahead of time. In the case of the risk neutral strategy its functional form is $s(t)=b_{i}^{k}t$ for firm $i$, type $k$. Therefore it is clear that the expected payoff function, see Eq. 2.13, for firm $i$, type $k$ is
$-\sum_{m=1}^{K_{i}}\pi_{i}^{k}(m)\int_{0}^{1}(\dot{s}_{i}^{k}+\dot{s}_{s_{\!\sm i}}^{m}+b_{i}^{k})\dot{s}_{i}^{k}+\kappa_{i}^{k}(s_{i}^{k}+s_{s_{\!\sm i}}^{m}+b_{i}^{k}\,t)\dot{s}_{i}^{k}\,\mathrm{dt}$ (3.38)
Now using the Euler-Lagrange equation we see that the strategy that minimizes the expected payoff for firm $i$ when its active type is $k$, is given by the following differential equation in terms of the $s_{s_{\!\sm i}}^{k}$:
$\ddot{s}_{i}^{k}+\frac{1}{2}\sum_{m=1}^{K_{i}}\ddot{s}_{s_{\!\sm i}}^{m}=-\frac{1}{2}\sum_{m=1}^{K_{i}}\pi_{i}^{k}(m)\kappa_{i}^{k}\dot{s}_{s_{\!\sm i}}^{m}-\frac{1}{2}\kappa_{i}^{k}b_{i}^{k}$ (3.39)
@@
$\kappa:=(\kappa_{1}^{1},\dots,\kappa_{1}^{K_{1}},\kappa_{2}^{1},\dots,\kappa_{2}^{K_{2}})^{\top}$ (3.41)
Now write * for element wise product and $v=\dot{s}$ as in Section 3.6. it is straightforward to see that analogous to Eq. 3.10 we have:
$A\dot{v}=-Bv-\frac{1}{2}\kappa*b$ (3.42)
@@
$\dot{v}=-(A^{T}A)^{\dagger}A^{T}Bv-\frac{1}{2}(A^{T}A)^{\dagger}A^{T}\kappa*b$ (3.43)
@@
$c:=-\frac{1}{2}(A^{T}A)^{\dagger}A^{T}\kappa*b$ (3.44)
@@
    $\begin{split}s_{1}^{*}&=(s_{1}^{1},\dots,s_{1}^{K})\\ s_{2}^{*}&=(s_{2}^{1},\dots,s_{1}^{M})\end{split}$ (3.49)
@@
    The vector $c$ is as in Eq. 3.44 and encapsulates the intensity of trading of the non-strategic firm in terms of the products of the market-impact parameters and the believed target sizes $\kappa_{i}^{k}b_{i}^{k}$.
@@
    One can see that when $b=0$ (that is $b_{i}^{k}=0$ for all $i,k$) then $\kappa*b=0$ and therefore $c=0$ and examining Eq. 3.47 shows $v_{0}=(e^{M}-I)^{-1}Mf=M(e^{M}-I)^{-1}f$.
In section
@@ 5 Two-firm equilibrium examples
Figure 1 displays the trading strategies for firms 1 and 2 in a 2x2 plot grid, each sub-plot of which shows one possible pair of active types for the firms, with the $(i,j)$ subplot showing the case where firm 1’s active type is $i$ and firm 2’s is $j$. To demonstrate the difference between complete and incomplete trading games, each plots also has an inset plot that shows what the strategies would be if the active types in the primary plot were the active types with certainty. In terms of Section 2.3 these plots show the normalized strategies (see Definition LABEL:def:normalized-strat) where $s_{1}^{*}=(s_{1}^{1},s_{1}^{2})$ consists of the firm 1 strategies (solid lines) for $t_{1}^{1}$ and $t_{1}^{2}$ in the top row of plots and $s_{2}^{*}$ consists of the firm 2 strategies (dash-dotted lines) in the first column.
@@ Discussion
Consider the following differences according to type pairs, referencing the figure.
@@
    When active types are $(1,1)$ (the top-left plot), firm 2 trades a mildly eager strategy (the dash-dotted line in the primary plot) while firm 1 trades a somewhat risk-averse strategy (the green line). Meanwhile, examining the inset which shows the strategies that would be traded if firm 1 were Type 1 and firm 2 were Type 2 with certainty.
@@
    When the active types are (2, 1) (the bottom-left plot), firm 2 again trades a modestly eager strategy while firm 1 trades an S-shape strategy that closely hugs the risk-neutral one. Meanwhile in the inset, where firm 1 is Type 2 and firm 2 is Type 1 with certainty firm 1 trades an eager strategy.
@@
    On the other hand, in the top-right plot when active types are (1, 2), the strategy shapes for firms 1 and 2 are more similar to the case when there is only one possible type per firm. Even in this case, however, the strategies are not the same. Both strategies in the latter case are more extreme in the sense that their farthest deviation from the risk-neutral strategy is greater then when their are two types per firm.
It can be challenging to understand why the shapes of normalized strategies are as they are. To begin, we reason through case where the active type of firm 1 and firm 2 is one as is displayed in the top-left plot of Figure 1 where firm 1’s strategy is risk-averse (convex) when there are two possible types but concave when there is one.
@@
    firm 2 is type 1 with a probability of 67%. In this case firm 2 believes $\kappa$ is 2.0 and has a target size of 7.0. In this case firm 2 trades a modestly eager strategy because it believes $\kappa$ is relatively small and it has a relatively large target size; or
@@
    firm 2 is type 2 with a probability of 33% and in this case firm 2 believes $\kappa$ is 15.0 with a target size of 5.0. With this very large $\kappa$ and a size at least as large as both of firm 1’s target sizes (firm 1’s target sizes are 3.0 and 5.0 for types 1 and 2 respectively), firm 2 trades a very eager strategy in order to get ahead of what it believes will be a significant amount of permanent impact.
@@
In order to demonstrate the impact of firm 1’s subjective probabilities on its strategy in Figure 2 we display firm 1 and firm 2’s strategy for precisely the same type data as in this section (and as displayed in Table 2) and with an almost identical common prior probability matrix as in Eq. 5.1 except that in each successive plot from left to right we decrease the probability that both firm 1 and firm 2 have active type 1.
@@
We now discuss implementation costs of the strategies in a little more detail. Consider when firm 1’s active type is 1. Then its normalized expected cost is 14.71 and in this case its believes that $\kappa=1.0$, meaning this is the value of the market impact parameter used in selecting its type 1 strategy. Meanwhile its actual cost depends on a few things, first among them is what the active type firm 2 is, which firm 1 does have ex ante knowledge of. In addition, it depends on the actual $\kappa$ that determines how the firms’ trading impacts market prices [^11] More generally the actual costs depend on the actual microstructure details governing how the market absorbs trades, and as discussed earlier one can make a more expansive Bayesian game that captures this uncertainty as well..
@@
Now referencing Table 2, we see that when firm 1’s active type is 1 then firm 1’s implementation costs are 47.4 and 37.5 when firm 2’s active type is 1 or 2 respectively. At first blush this seems odd because consulting Figure 1, when firm 2’s active type is 2 it trades considerably more aggressively through time $t=0.6$ (that is, through 60% of the total time trading). Should that not push prices up so that when firm 1 as type 1 faces firm 2 as type 2 it pays more than when firm 2 is type 1?
@@
Figure 4: Cumulative costs for the trading strategies depicted in Figure 1 with total implementation costs as in Table 2 for the evaluation $\kappa$ set to 1.0.

### v2

@@ 3.6 Bayesian equilibrium
Therefore in the case where $A$ is singular (or could be singular) we can use the alternative definition of $M$ given my
$M:=-(A^{T}A)^{-1}A^{T}B$ (3.23)
In the case $A$ is non-singular this is equivalent to 3.19 because as is well-known $(A^{T}A)^{-1}A^{T}=A^{-1}$ (when $A$ is non-singular ) [^14] This is because when $A$ is non-singular then $(A^{T}A)^{-1}=A^{-1}(A^{T})^{-1}$ then $(A^{T}A)^{-1}A^{T}=A^{-1}(A^{T})^{-1}A^{T}=A^{-1}$.. In the next section we use this to explicitly solve for the relevant strategies. In general we will proceed without reference to the normal-form equations but note that these should be employed when $A$ may not singular or nearly singular.
@@
We set our aim to find the most straightforward description of the matrix $C=-A^{-1}B$ as possible. To begin we construct matrices $\Pi_{1}$, $\Pi_{2}$ from the sub-blocks of equations (3.12) and (3.13) consisting of conditional probabilities:
@@
$C=-A^{-1}B$ $=\begin{pmatrix}-(I-\Pi_{1}\Pi_{2})^{-1}&\Pi_{1}(I-\Pi_{2}\Pi_{1})^{-1}\\ \Pi_{2}(I-\Pi_{1}\Pi_{2})^{-1}&-(I-\Pi_{2}\Pi_{1})^{-1}\end{pmatrix}\begin{pmatrix}0&D_{1}\Pi_{1}\\ D_{2}\Pi_{2}&0\end{pmatrix}$
@@
Note that Eq. 3.34 provides a complete description of a two-firm Bayesian equilibrium of the two specific strategies: $s_{1}^{*}=(s_{1}^{1},\dots,s_{1}^{K})$ and $s_{2}^{*}=(s_{2}^{1},\dots,s_{2}^{M})$ satisfying the boundary conditions in Eq. 3.32. We also note that from Eq. 4.2b we immediately have:
@@ 3.7 Two firms competing in the presence a non-strategic firm
To illustrate the tremendous generality and flexibility of types, we consider the following situation:
  1. 1.
    There are two firms in competition with one another.
  2. 2.
    There is the possibility of a third, non-strategic firm trading the same stock as firms 1 and 2 over the same period of time. The non-strategic firm is still in competition with the other two firms but employs none of the game-theoretic logic of strategy selection.
  3. 3.
    We call the first two firms the strategic firms and the add to their type data a third item which describes their belief concerning the non-strategic firms target size (see Section 2.3).
  4. 4.
    Thus the type data for each firm in this situation is a triple consisting of the each types belief in what the market impact coefficient is, what their target size is and what the non-strategic firm’s target size is.
Below we will analyze how to apply the machinery of Bayesian games to this situation. First, however, we show that the above setup is in some sense a natural generalization of the Almgren-Chriss model.
@@
    We will write $K$ and $M$ for the number of types for firm 1 and 2 respectively.
@@
    Each possible active type of firms 1 and 2 believes that the non-strategic firm will trade a risk-neutral strategy with a specific target size. We write $b_{i}^{k}$ for firm $i$, type $k$’s belief about the target size of the non-strategic firm. Note that it is possible that some of the $b_{i}^{k}$ are equal to zero meaning that firm $i$, type $k$ believes there not a non-strategic investor.
@@
    The type data for firm $i$ when its active type is $k$ is a triple $(\kappa_{i}^{k},f_{i}^{k},b_{i}^{k})$, where $\kappa_{i}^{k}$ is firm $i$’s market impact coefficient, $f_{i}^{k}$ is its target size and $b_{i}^{k}$ represents the target size that firm $i$ (when its active type is $k$) believes the non-strategic firm will trade.
@@
    There is a common prior probability matrix $P$, see Eq. 2.3, such that $p_{km}$ is the joint probability that firm 1’s active type is $k$ and firm 2’s is $m$. Note that unlike before, each firm’s type includes their view of the non-strategic firm’s target size.
@@
We note that it is possible and straightforward to extend the analysis we do so that some firms types believe that the non-strategic firm trades some other strategy other than the risk-neutral strategy and our analysis directly extends to this case so long as the other other strategy’s functional form is fixed and known ahead of time. In the case of the risk neutral strategy its functional form is $s(t)=b_{i}^{k}t$ for firm $i$, type $k$. Therefore it is clear that the expected payoff function, see Eq. 2.14, for firms $1$ and $2$ when their active types are $k$ and $m$ respectively are given by:
Firm 1 $=-\sum_{m=1}^{M}\pi_{1}^{k}(m)\int_{0}^{1}(\dot{s}_{1}^{k}+\dot{s}_{2}^{m}+b_{1}^{k})\dot{s}_{1}^{k}+\kappa_{i}^{k}(s_{1}^{k}+s_{2}^{m}+b_{1}^{k}\,t)\dot{s}_{1}^{k}\,\mathrm{dt}$ (3.39)
Firm 2 $=-\sum_{k=1}^{K}\pi_{2}^{m}(k)\int_{0}^{1}(\dot{s}_{1}^{k}+\dot{s}_{2}^{m}+b_{2}^{k})\dot{s}_{2}^{m}+\kappa_{i}^{k}(s_{1}^{k}+s_{2}^{m}+b_{1}^{k}\,t)\dot{s}_{2}^{m}\,\mathrm{dt}$ (3.40)
Now using the Euler-Lagrange equation we see that the strategy that minimizes the expected payoff for firm $1$ (firm 2 is analogous) when its active type is $k$, is given by the following differential equation in terms of the $s_{s_{\!\sm i}}^{k}$:
$\ddot{s}_{1}^{k}+\frac{1}{2}\sum_{m=1}^{M}\ddot{s}_{2}^{m}=-\frac{1}{2}\sum_{m=1}^{M}\pi_{1}^{k}(m)\kappa_{1}^{k}\dot{s}_{2}^{m}-\frac{1}{2}\kappa_{i}^{k}b_{1}^{k}$ (3.41)
@@
$\kappa:=(\kappa_{1}^{1},\dots,\kappa_{1}^{K},\kappa_{2}^{1},\dots,\kappa_{2}^{M})^{\top}$ (3.43)
Now write $\star$ for element wise product and $v=\dot{s}$ as in Section 3.6. it is straightforward to see that analogous to Eq. 3.10 we have:
$A\dot{v}=-Bv-\frac{1}{2}\kappa\star b$ (3.44)
@@
$\dot{v}=-(A^{T}A)^{\dagger}A^{T}Bv-\frac{1}{2}(A^{T}A)^{\dagger}A^{T}\kappa\star b$ (3.45)
@@
$c:=-\frac{1}{2}(A^{T}A)^{\dagger}A^{T}\kappa\star b$ (3.46)
@@
    $\begin{split}s_{1}^{*}&=(s_{1}^{1},\dots,s_{1}^{K})\\ s_{2}^{*}&=(s_{2}^{1},\dots,s_{2}^{M})\end{split}$ (3.51)
@@
    The vector $c$ is as in Eq. 3.46 and encapsulates the intensity of trading of the non-strategic firm in terms of the products of the market-impact parameters and the believed target sizes $\kappa_{i}^{k},b_{i}^{k}$.
@@
    One can see that when $b=0$ (that is $b_{i}^{k}=0$ for all $i,k$) then $\kappa*b=0$ and therefore $c=0$. Examining Eq. 3.49 this shows $v_{0}=(e^{M}-I)^{-1}Mf=M(e^{M}-I)^{-1}f$.
@@ 5 Two-firm equilibrium examples
Figure 1 displays the trading strategies for firms 1 and 2 in a plot, each sub-plot of which shows one possible pair of active types for the firms, with the $(i,j)$ subplot showing the case where firm 1’s active type is $i$ and firm 2’s is $j$. To demonstrate the difference between complete and incomplete trading in competition, each plot also displays an inset that shows what the strategies would be if the active types in the primary plot were the active types with certainty. In terms of Section 2.3 these plots show the strategy profiles (see Definition 2.5) where $s_{1}^{*}=(s_{1}^{1},s_{1}^{2})$ consists of the firm 1 strategies (solid lines) for $t_{1}^{1}$ and $t_{1}^{2}$ in the top row of plots and $s_{2}^{*}$ consists of the firm 2 strategies (dash-dotted lines) in the first column.
@@ Discussion
Consider the following differences according to type pairs, referencing the figure. We will write $(k,m)$ for the tuple representing firm 1 when its active type is $k$ and firm 2 when its active type is simultaneously $m$.
@@
    When the active types of firms 1 and 2 are $(1,1)$ (the leftmost plot in Figure 1), firm 2 trades a mildly eager strategy (the dash-dotted line in the primary plot) while firm 1 trades a somewhat risk-averse strategy (the green line). Meanwhile, examining the inset which shows the strategies that would be traded if firm 1 were Type 1 and firm 2 were Type 2 with certainty.
@@
    When the active types are (2, 1) (the third plot in Figure 1), as with the case of active types (1, 1) firm 2 again trades a modestly eager strategy while firm 1 trades an S-shape strategy that closely hugs the risk-neutral one. Meanwhile in the inset, where firm 1 is Type 2 and firm 2 is Type 1 with certainty firm 1 trades an eager strategy.
@@
    On the other hand, when active types are (1, 2) (the second plot in Figure 1), the strategy shapes for firms 1 and 2 are more similar to the case when there is only one possible type per firm. Even in this case, however, the strategies are not precisely the same as in the complete information case. Both strategies in the latter case are more extreme in the sense that their farthest deviation from the risk-neutral strategy is greater then when their are two types per firm.
  4. 4.
    Finally when both types are (2, 2) (the rightmost plot in Figure 1) provides an interesting example of the effect that uncertainty has on trading strategies. Note that in this case firm 2 believes the market impact coefficient is 15.0 and has a target size of 5.0. So why does it trade so aggressively? This seems substantially different than the complete information case of [Chr24b]. The main reason is because firm 2 knows that firm 1 believes the market impact coefficient is either 1 or 3 and therefore is not concerned with the market impact firm 2 will cause. In the complete information case in [Chr24b] firms share a common believe about the market impact parameter.
It can be challenging to understand why the shapes of optimal strategies are as they are. To begin, we reason through case where the active types of firm 1 and firm 2 are both one as displayed in the leftmost plot of Figure 1 where firm 1’s strategy is risk-averse (convex) when there are two possible types but concave when there is one.
@@
    Firm 2’s active type is 1 with a probability of 67%, believes $\kappa$ is 2.0 and has a target size of 7.0. In this case firm 2 trades a modestly eager strategy because it believes $\kappa$ is relatively small and it has a relatively large target size; or
@@
    Firm 2’s active type is 2 with a probability of 33%, believes $\kappa$ is 15.0 and has a target size of 5.0. With the very large $\kappa$ of 15.0 and a target size at least as large as both of firm 1’s active types’ target sizes (firm 1’s target sizes are 3.0 and 5.0 for types 1 and 2 respectively), firm 2 trades a very eager strategy in order to get ahead of what it believes will be a significant amount of permanent impact.
@@
In order to demonstrate the impact of firm 1’s subjective probabilities on its strategy in Figure 2 we display firm 1 and firm 2’s strategy for precisely the same type data as already displayed in Figure 1 (and as in Table 3) and with an almost identical common prior probability matrix as in Eq. 5.1 except that in each successive plot from left to right we decrease the probability in the common prior that both firm 1 and firm 2 have active type 1.
We display result of this in the plot titles by showing the conditional probabilities. For example, firm 1’s conditional probability of firm 2 being type 1 from left to right are given by $\pi_{1}^{1}(1)=67\%$, $\pi_{1}^{1}(1)=45\%$, $\pi_{1}^{1}(1)=23\%$ and $\pi_{1}^{1}(1)=2\%$. Thus in each successive plot firm 1 wen its active type is 1 holds a subjective probability that firm 2 is type 1 that decreases from 67% to 2%. Meanwhile firm 2, when its active type is 1, has a subjective belief concerning firm 1’s active type being 1 that drops from 73% to 6%. The impact on trading is mainly that firm 1 trades increasingly slowly. Let’s understand why. Consider the rightmost plot where firm 1’s active type is 1 and therefore it believes the market impact coefficient is small at $\kappa=1.0$ and further it believes with close to certainty (94%) that firm 2’s active type is 2 and thus its target size is 5.0 (as opposed to 7.0) with close to certainty. Thus firm 1’s beliefs are that $\kappa$ is small and that firm 2 is very likely trading the smaller of its two possible sizes. Moreover, firm 1 believes that firm 2 that the market impact coefficient is very large (because its active type is very likely 2 firm 2 believes $\kappa=15.0$) . With a $\kappa$ that large, firm 1 believes that firm 2 will trade eagerly and can benefit from the late selling that firm 2 must do to hit its target size.
@@
Given the strategies above, we discuss their implementation costs as described in Section 4. We continue to use the type data from Table 3 which we repeat here for convenience.
Type 1 Type 2
Firm 1 $t_{1}^{1}=(1.0,3.0)$ $t_{1}^{2}=(3.0,5.0)$
Firm 2 $t_{2}^{1}=(2.0,7.0)$ $t_{2}^{2}=(15.0,5.0)$
Each type $t_{i}^{k}=(\kappa_{i}^{k},f_{i}^{k})$ consists of a market impact parameter $\kappa_{i}^{k}$ and a target size $f_{i}^{k}$. For example, firm 1’s type 1 has $\kappa_{1}^{1}=1.0$ and $f_{1}^{1}=3.0$. In Table 3 we display the costs and expected costs as follows.
@@
Firm 1 Firm 2 Firm 1 Firm 2
Type $\kappa$ $f$ Type $\kappa$ $f$  Eval $\kappa_{e}$ Cost Exp. Cost Exp.
$t_{1}^{1}$ 1.0 3.0 $t_{2}^{1}$ 2.0 7.0 1.0 47.4 44.1 102.7 153.2
1.0 3.0 2.0 7.0 2.0 65.2 135.0
1.0 3.0 2.0 7.0 3.0 83.0 167.2
1.0 3.0 2.0 7.0 15.0 296.0 554.1
$t_{1}^{1}$ 1.0 3.0 $t_{2}^{2}$ 15.0 5.0 1.0 37.5 44.1 113.2 285.5
1.0 3.0 15.0 5.0 2.0 57.4 125.3
1.0 3.0 15.0 5.0 3.0 77.3 137.4
1.0 3.0 15.0 5.0 15.0 316.1 282.4
$t_{1}^{2}$ 3.0 5.0 $t_{2}^{1}$ 2.0 7.0 1.0 93.1 156.7 124.9 153.2
3.0 5.0 2.0 7.0 2.0 125.1 164.9
3.0 5.0 2.0 7.0 3.0 157.1 204.9
3.0 5.0 2.0 7.0 15.0 541.1 684.9
$t_{1}^{2}$ 3.0 5.0 $t_{2}^{2}$ 15.0 5.0 1.0 84.9 156.7 129.9 285.5
3.0 5.0 15.0 5.0 2.0 120.7 144.1
3.0 5.0 15.0 5.0 3.0 156.4 158.3
3.0 5.0 15.0 5.0 15.0 585.7 329.0
Table 3: Costs for Firms 1 and 2 for the case of two firms and 2 types. The costs displayed have not been normalized.
We repeat the Table 2 but this time with the costs not normalized. We now discuss implementation costs of the strategies in a little more detail. Consider when firm 1’s active type is 1. Then its normalized expected cost is 14.71 and in this case its believes that $\kappa=1.0$, meaning this is the value of the market impact parameter used in selecting its type 1 strategy. Meanwhile its actual cost depends on a few things, first among them is what the active type firm 2 is, which firm 1 does have ex ante knowledge of. In addition, it depends on the actual $\kappa$ that determines how the firms’ trading impacts market prices [^15] More generally the actual costs depend on the actual microstructure details governing how the market absorbs trades, and as discussed earlier one can make a more expansive Bayesian game that captures this uncertainty as well..
@@
Now referencing Table 3, we see that when firm 1’s active type is 1 then firm 1’s implementation costs are 15.81 and 14.68 when firm 2’s active type is 1 or 2 respectively. At first blush this seems odd because consulting Figure 1, when firm 2’s active type is 2 it trades considerably more aggressively up to time $t=0.6$ (that is, through 60% of the total time trading). Should that not push prices up so that when firm 1 as type 1 faces firm 2 as type 2 it pays more than when firm 2 is type 1?
@@
Figure 4: Cumulative costs for the trading strategies depicted in Figure 1 with total implementation costs as in Table 3 for the evaluation $\kappa$ set to 1.0. One can compare these costs to the implementation costs in Table 3.